A +40uC charged is placed at the origin. Calculate the magnitude and direction of the field created by the point charged at the following A point (5m,0) B. (3m,4m)

1
Expert's answer
2022-02-21T14:59:54-0500

Part A


"E=\frac{kq}{r^2}\\r=5m"


"\therefore\>E=\frac{9×10^9×40×10^{-6}}{5^2}"

"=1.44\u00d710^4N\/C"


Toward positive x-axis


Part B


"r=\sqrt{3^2+4^2}=5m"


"|E|=1.44\u00d710^4N\/C"


"\theta=" tan"^{-1}\>\frac{4}{3}=53.13"


"E=1.44\u00d710^4N\/C"

Along a line 53.13° to x-axis directed away from origin.